The pressure in a constant-volume gas thermometer is 0.700 atm at 100°C and 0.512 atm at 0°C. (a) What is the temperature when the pressure is 0.0738 atm? °C (b) What is the pressure at 358°C?
Answer #1

Volume of the gas is constant

Pressure P1 = 0.7 atm at temperature T1 = 100C°

Pressure P2 = 0.512 atm at T2 = 0c°

1atm = 101325 pa

1C = 274.15 Kelvin

By the ideal gas law

PV = nRT

Where R = gas constant

n = number of molecules

V= volume

Hence

P directly proportional to T

P1= 0.7atm = 0.7* 101325 = 70927.5 pa

T1= 100c = ( 100+274.15)= 374.15 Kelvin

P2=0.0738 atm = 0.0738*101325=7477.785 pa

T2= 39.45 Kelvin

T2= ( 39.45 - 274.15) = -234.7 c

This is the answer for a

b) let the pressure at T3= 358C° is P3

T3= (358 +274.15) Kelvin

T3= 632.15 Kelvin

P3= 119836.48

P3= 1.1827 atm

This is the pressure at 358°C